# Definability of mad families of vector spaces and two local Ramsey theories

### v4

@@ Abstract.
Let $E$ be a vector space over a countable field of dimension $\aleph_{0}$. Two infinite-dimensional subspaces $V,W\subseteq E$ are almost disjoint if $V\cap W$ is finite-dimensional. This paper provides some improvements on results about the definability of maximal almost disjoint families (mad families) of subspaces in [S19]. We construct a full mad family of block subspaces in $\mathsf{ZFC}$, answering a problem by Smythe in the positive. A variant of this construction shows that there exists a completely separable mad family of block subspaces in $\mathsf{ZFC}$. We also discuss the abstract Mathias forcing introduced by Di Prisco-Mijares-Nieto in [DMN15], and apply it to show that in the Solovay’s model obtained by the collapse of a Mahlo cardinal, there are no full mad families of block subspaces over $\mathbb{F}_{2}$.
@@ 1. Introduction
Let $E$ be a vector space over a countable field $\mathbb{F}$ of dimension $\aleph_{0}$. In [S19], Smythe studied almost disjoint families infinite-dimensional subspaces $V,W\subseteq E$, i.e. families of infinite-dimensional subspaces of $E$ which pairwise intersections are finite-dimensional. Our paper focuses on the following problems raised by Smythe in [S19].
@@
Horowitz-Shelah proved in [HS22] that there are no analytic mad families of subspaces over $\mathbb{F}_{2}$, but the first problem remains largely open for other fields. Smythe remarked in [S19] that if there is a supercompact cardinal, then every set in Solovay’s model is $\mathcal{H}$-strategically Ramsey, so there are no full mad families (the definition of a full mad family will be introduced in the subsequent paragraph). Strategically Ramsey subsets of $E^{[\infty]}$ were introduced in [R08] and [R09] to study the combinatorics of Banach space theory and dichotomies, and Smythe introduced a localised version of the strategically Ramsey property in [S18]. To the author’s knowledge, the precise large cardinal strength needed for the non-existence of mad families in $\mathbf{L}(\mathbb{R})$ is mostly unknown.
Inspired by Mathias’ original proof of the non-existence of analytic mad families of $\omega]^{\omega}$, Smythe approached this problem using a local Ramsey theoretic approach in [S19], replacing $\mathcal{H}$-Ramsey with $\mathcal{H}$-strategically Ramsey. Given an almost disjoint family $\mathcal{A}$ of vector spaces, he extended Mathias’ definition of the “happy family” $\mathfrak{F}_{\mathcal{A}}$ of subsets of $\omega$, to a “family” $\mathcal{H}(\mathcal{A})$ of subspaces of $E$. His approach requires $\mathcal{H}(\mathcal{A})$ to be full, a property which closely resembles the pigeonhole property of coideals of $\omega]^{\omega}$. Calling a mad family $\mathcal{A}$ full if $\mathcal{H}(\mathcal{A})$ is full, Smythe proved the following theorem:
Theorem 1.2 (Smythe [S19]).
There exist no analytic full mad families of block subspaces.
Block subspaces are subspaces with a basis of vectors with strictly increasing support (see §2 for a precise definition). Block subspaces are more combinatorially tame, and since every infinite-dimensional subspace contains an infinite-dimensional block subspace, an almost disjoint family of subspaces is maximal with respect to block subspaces iff it is maximal.
While it is clear from Smythe’s proof that the fullness property plays an important role in understanding the definability of mad families of subspaces, the extent which mad families of subspaces satisfy such property remains poorly understood. In particular, Smythe asked the following question in [S19]:
Problem 1.3 (Smythe, [S19]).
@@
Another obstacle to understanding the definability of mad families of subspaces is the restriction to mad families of block subspaces. It is unknown if Theorem 1.2 holds for arbitrary mad families of subspaces (except for when $|\mathbb{F}|=2$). Furthermore, consider the following two cardinals:
@@
Our method is inspired by a variant of the construction of a completely separable mad family of $\omega]^{\omega}$ in [MRS14] assuming $\mathfrak{s}\leq\mathfrak{a}$, provided by Guzmán in [G21]. Construction of a completely separable mad family of $\omega]^{\omega}$ has been extensively studied in the past two decades. Shelah attained a breakthrough in [Sh11], where he showed that a completely separable mad family assuming either $\mathfrak{c}<\aleph_{\omega}$, or there is no inner model with a measurable cardinal. However, the existence of a completely separable mad family in $\mathsf{ZFC}$ alone remains open. We shall show that the natural generalisation of a completely separable mad family to vector spaces (see Definition 5.6) exists in $\mathsf{ZFC}$.
@@
Our constructions for Theorem 1.5 and 1.6 utilises various results from the local Ramsey theory of block subspaces developed by Smythe in [S18] and [S19], along with a few additional results that we shall prove.
The second part in this paper addresses the existence of full mad families of block subspaces over $\mathbb{F}_{2}$ in Solovay’s model. We utilise the abstract Mathias forcing developed by Di Prisco-Mijares-Nieto in [DMN15] - the abstract Mathias forcing was developed to expand the local Ramsey theory of $\omega]^{\omega}$ to topological Ramsey spaces (i.e. closed triples $(\mathcal{R},\leq,r)$ satisfying the axioms A1-A4 introduced by Todorčević in [T10]). For example, if $E$ is a countable vector space over $\mathbb{F}_{2}$, then the pigeonhole principle holds, so it is a topological Ramsey space.
Let $\operatorname{Coll}(\omega,<\!\kappa)$ be the Lévy collapse, which collapses the uncountable cardinal $\kappa$ to $\omega_{1}$. Using the abstract Mathias forcing, we may apply the abstract Mathias forcing to obtain the following variant of a celebrated theorem of Mathias:
@@
Let $\kappa$ be a Mahlo cardinal, and let $G$ be a $\operatorname{Coll}(\omega,<\!\kappa)$-generic filter. Then there are no full mad families of block subspaces over $\mathbb{F}_{2}$ in $\mathbf{L}(\mathbb{R})^{\mathbf{V}[G]}$.
@@
    We provide a concise introduction to various definitions and notations to block subspaces of $E$.
@@
    We introduce some definitions and basic results about almost disjoint families of (block) subspaces. We show that $\mathfrak{a}_{\mathrm{vec},\mathbb{F}}^{*}$ is uncountable, addressing a remark by Smythe in [S19]. We also show that the semicoideal $\mathcal{H}^{-}(\mathcal{A})$ is selective for any almost disjoint family $\mathcal{A}$ - the non-requirement for $\mathcal{A}$ to be maximal is crucial in the proof of Theorem 1.5 and 1.6.
@@
    We apply various lemmas and results proven in §3 and §4 to provide a detailed proof of Theorem 1.5 and 1.6.
@@
    We provide an overview of the axioms of topological Ramsey spaces A1-A4 introduced by Todorčević, and the abstract local Ramsey theory developed in [DMN15]. We show that given a closed triple $(\mathcal{R},\leq,r)$ satisfying A1-A4, if $\mathcal{H}\subseteq\mathcal{R}$ is semiselective then $\mathbb{M}_{\mathcal{H}}$ satisfies the Mathias property and the Prikry property. This was proven in [DMN15], but they assumed that the almost reduction relation $\leq^{*}$ is transitive (which is not generally true, such as for the topological Ramsey space of strong subtrees). We provide a proof of this result without this assumption.
@@ 2. Vector spaces and block subspaces
We introduce some basic definitions and properties of vector spaces and block subspaces.
@@ 2.1. Setup and notations
Let $E$ be a vector space over a countable field $\mathbb{F}$, with a countable Hamel basis $(e_{n})_{n<\omega}$. For each $x\in E$, we may write $x:=\sum_{n<\omega}\lambda_{n}(x)e_{n}$, where $\lambda_{n}\in\mathbb{F}$. This allows us to define the support of a vector $x\in E$ by:
@@
This, in turn, gives us a partial order << on $E$, where for any two non-zero vectors $x,y\in E$:
@@
Since such a block basis is unique up to scaling, we may conflate block subspaces of $E$ with infinite block sequences.
@@
  • If $A=(x_{n})_{n<\omega}\in E^{[\infty]}$ and $N<\omega$, we write $A/N:=(x_{n})_{n\geq N}$.
@@
  • If $A\in E^{[<\infty]}$ and $a\in E^{[<\infty]}$ we write $d_{A}(a):=\min\{N<\omega:\max(\operatorname{supp}(a))<A/N\}$.
@@
    $\langle(x_{n})_{n<\alpha}\rangle:=\operatorname{span}\{x_{n}:n<\alpha\}.$
@@ 2.2. Intersection of block subspaces
Given $A,B\in E^{[\infty]}$, if $\langle A\rangle\cap\langle B\rangle$ is infinite-dimensional, then by the remark above it contains a block subspace. The following lemma shows that $\langle A\rangle\cap\langle B\rangle$ is itself a block subspace. This observation is central in ensuring that the proof of Theorem 1.5 and 1.6 do not require additional set-theoretic hypotheses.
@@
For an arbitrary $v\in E$ and $k\in\operatorname{supp}(v)$, we also write $v(k):=e_{k}^{*}(v)\in\mathbb{F}$ to be the unique coefficient such that $v=u+v(k)e_{k}+w$ for some vectors $u<e_{k}<w$ (or $u=0$ or $w=0$).
@@
$v_{0}$ $=\sum_{i\in\operatorname{supp}_{A}(v_{0})}\lambda_{i}^{0}x_{i}=\sum_{j\in\operatorname{supp}_{B}(v_{0})}\mu_{j}^{0}y_{j},$ (2.1)
$v_{1}$ $=\sum_{i\in\operatorname{supp}_{A}(v_{1})}\lambda_{i}^{1}x_{i}=\sum_{j\in\operatorname{supp}_{B}(v_{1})}\mu_{j}^{1}y_{j}.$ (2.2)
where $\lambda_{i}^{0},\lambda_{i}^{1},\lambda_{j}^{0},\lambda_{j}^{1}$ are all non-zero. We let:
$w:=\sum_{i\in\operatorname{supp}_{A}(v_{0})\cap\operatorname{supp}_{A}(v_{1})}\lambda_{i}^{0}x_{i},$
@@
By the symmetry of the argument, it suffices to show that $\operatorname{supp}(w)\subseteq\operatorname{supp}(w^{\prime})$, and $w(k)=w^{\prime}(k)$ for all $k\in\operatorname{supp}(w)$. Let $k\in\operatorname{supp}(w)$, so $k\in\operatorname{supp}(x_{i})$ for some $i\in\operatorname{supp}_{A}(v_{0})\cap\operatorname{supp}_{A}(v_{1})$. By the second inequalities in (2.1) and (2.2), we have that $k\in\operatorname{supp}(y_{j})$ for some $j\in\operatorname{supp}_{B}(v_{0})$, and $k\in\operatorname{supp}(y_{j^{\prime}})$ for some $j^{\prime}\in\operatorname{supp}_{B}(v_{1})$. Since $B$ is a block sequence, $k\in\operatorname{supp}(y_{j})\cap\operatorname{supp}(y_{j^{\prime}})$ implies that $j=j^{\prime}$. Therefore, $j\in\operatorname{supp}_{B}(v_{0})\cap\operatorname{supp}_{B}(v_{1})$, so $k\in\operatorname{supp}(w^{\prime})$. By (2.1), we can conclude that:
$w(k)=\lambda_{i}^{0}x_{i}(k)=v_{0}(k)=\mu_{j}^{0}y_{j}(k)=w^{\prime}(k),$
@@
$z_{m}$ $=\sum_{i\in\operatorname{supp}_{A}(z_{m})}\lambda_{i}x_{i}=\sum_{j\in\operatorname{supp}_{B}(z_{m})}\mu_{j}y_{j},$
where $\lambda_{i},\mu_{j}$ are non-zero. Let $i_{0}:=\max\{i\in\operatorname{supp}_{A}(z_{m}):i<n\}$ and $i_{1}:=\min\{i\in\operatorname{supp}_{A}(z_{m}):i>n\}$, and let $k_{0}:=\max(\operatorname{supp}(x_{i_{0}}))$ and $k_{1}:=\min(\operatorname{supp}(x_{i_{1}}))$. We define:
$w:=\sum_{\begin{subarray}{c}i\in\operatorname{supp}_{A}(z_{m})\\ \max(\operatorname{supp}(x_{i}))\leq k_{0}\end{subarray}}\lambda_{i}^{m}x_{i}.$
@@
Observe that for each $j\in\operatorname{supp}_{B}(z_{m})$, either $k_{0}<\min(\operatorname{supp}(y_{j}))$ or $\max(\operatorname{supp}(y_{j}))<k_{1}$ – otherwise, if $j\in\operatorname{supp}_{B}(z_{m})$ is such that $\min(\operatorname{supp}(y_{j}))\leq k_{0}<k_{1}\leq\max(\operatorname{supp}(y_{j}))$, then by the block sequence property of $B$, we have that for all $j^{\prime}\neq j$, either $\max(\operatorname{supp}(y_{j^{\prime}}))<k_{0}$ or $k_{1}<\min(\operatorname{supp}(y_{j^{\prime}}))$. Since $k_{0}<\min(\operatorname{supp}(x_{n}))\leq\max(\operatorname{supp}(x_{n}))<k_{1}$ and $j\notin\operatorname{supp}_{B}(z_{n})$, this implies that:
$\operatorname{supp}(x_{n})\cap\bigcup_{j\in\operatorname{supp}_{B}(z_{n})}\operatorname{supp}(y_{j})=\emptyset,$
@@
Note that for all $k\in\operatorname{supp}(z_{m})$, either $k\leq k_{0}$ or $k_{1}\leq k$. Thus, for all $j\in\operatorname{supp}_{B}(z_{m})$, since $\operatorname{supp}(y_{j})\subseteq\operatorname{supp}(z_{m})$, if $\max(\operatorname{supp}(y_{j}))<k_{1}$ then $\max(\operatorname{supp}(y_{j}))\leq k_{0}$. Similarly, if $k_{0}<\min(\operatorname{supp}(y_{j}))$ then $k_{1}\leq\min(\operatorname{supp}(y_{j}))$. Therefore we have that:
@@
then we have that $\operatorname{supp}(w)=\operatorname{supp}(w^{\prime})$. Furthermore, for any $k\in\operatorname{supp}(w)$, $w(k)=z_{m}(k)=w^{\prime}(k)$, so $w=w^{\prime}$. It’s clear from the definition that $w\in\langle A\rangle$ and $w^{\prime}\in\langle B\rangle$, so we have that $w\in\langle A\rangle\cap\langle B\rangle$, as desired. ∎

### v5

# Full mad families of vector spaces and two local Ramsey theories

@@ Abstract.
Let $E$ be a vector space over a countable field of dimension $\aleph_{0}$. Two infinite-dimensional subspaces $V,W\subseteq E$ are almost disjoint if $V\cap W$ is finite-dimensional. This paper provides some improvements on results about the definability of maximal almost disjoint families (mad families) of subspaces in [24]. We construct a full mad family of block subspaces in $\mathsf{ZFC}$, answering a problem by Smythe in the positive. A variant of this construction shows that there exists a completely separable mad family of block subspaces in $\mathsf{ZFC}$. We also discuss the abstract Mathias forcing introduced by Di Prisco-Mijares-Nieto in [6], and apply it to show that in the Solovay’s model obtained by the collapse of a Mahlo cardinal, there are no full mad families of subspaces over $\mathbb{F}_{2}$.
@@ 1. Introduction
Let $E$ be a vector space over a countable field $\mathbb{F}$ of dimension $\aleph_{0}$. In [24], Smythe studied almost disjoint families of infinite-dimensional subspaces $V,W\subseteq E$, i.e. families of infinite-dimensional subspaces of $E$ whose pairwise intersections are finite-dimensional, and raised the following questions.
@@
The study of almost disjoint families in various linear algebraic notions, or applications of the theory of almost disjoint families to linear algebra or functional analysis, predates Smythe’s work in [24]. In [1], Baumgartner-Spinas studied the existence of quadratic spaces, and used almost disjoint families in $\omega]^{\omega}$ to show that under the assumption $\mathfrak{b}=\aleph_{1}$, there exists an $\aleph_{1}$-dimensional (**)-space over a class of countable fields. Possible cardinalities of variants of almost disjoint families, including almost orthogonal families, of infinite-dimensional subspaces of a separable Hilbert space were also studied in [2] and [30]. Kolman also studied almost disjoint families of infinite-dimensional subspaces with no additional topological assumptions in [12] (which he called almost disjoint packings), but he assumed that almost disjoint families satisfy a property stronger than pairwise almost disjointness (see also footnote 2 of [24]).
Horowitz-Shelah proved in [10] that there are no analytic mad families of subspaces over $\mathbb{F}_{2}$, but the first problem remains open for other fields. Smythe remarked in [24] that if there is a supercompact cardinal, then every set in $\mathbf{L}(\mathbb{R})$ is $\mathcal{H}$-strategically Ramsey, so there are no full mad families of subspaces there (the definition of a full mad family of subspaces will be introduced in the subsequent paragraph). Strategically Ramsey subsets of $E^{[\infty]}$ were introduced in [18] and [19] to study the combinatorics of Banach space theory and dichotomies, and Smythe introduced a localised version of the strategically Ramsey property in [23]. To the author’s knowledge, the precise large cardinal strength needed for the non-existence of mad families of subspaces in $\mathbf{L}(\mathbb{R})$ is mostly unknown.
Inspired by Mathias’ original proof of the non-existence of analytic mad families in $\omega]^{\omega}$ in [13], Smythe approached this problem using a local Ramsey theoretic approach in [24], replacing $\mathcal{H}$-Ramsey with $\mathcal{H}$-strategically Ramsey. Given an almost disjoint family $\mathcal{A}$ of vector spaces, he extended Mathias’ definition of the “happy family” $\mathfrak{F}_{\mathcal{A}}$ of subsets of $\omega$, to a “family” $\mathcal{H}(\mathcal{A})$ of subspaces of $E$. His approach requires $\mathcal{H}(\mathcal{A})$ to be full, a property which closely resembles the pigeonhole property of coideals of $\omega]^{\omega}$. Calling a mad family $\mathcal{A}$ full if $\mathcal{H}(\mathcal{A})$ is full, Smythe proved the following theorem:
Theorem 1.2 (Smythe [24]).
There exist no analytic full mad families of subspaces.
Smythe proved this theorem by studying the combinatorial properties of families of block subspaces. Block subspaces are subspaces with a basis of vectors with strictly increasing support (see §2 for a precise definition). Block subspaces are more combinatorially tame, and since every infinite-dimensional subspace contains an infinite-dimensional block subspace, an almost disjoint family of subspaces is maximal with respect to block subspaces iff it is maximal.
While it is clear from Smythe’s proof that the fullness property plays an important role in understanding the definability of mad families of subspaces, the extent which mad families of subspaces satisfy this property remains poorly understood. In particular, Smythe asked the following question in [24]:
Problem 1.3 (Smythe [24]).
@@
Since every infinite-dimensional subspace contains an infinite-dimensional block subspace, not all generality is lost when restricting our attention to block subspaces. However, several open problems concerning differences between mad families of subspaces and of block subspaces remain. Consider the following two cardinals:
@@
Our method is inspired by a variant of the construction of a completely separable mad family in $\omega]^{\omega}$ in [14] assuming $\mathfrak{s}\leq\mathfrak{a}$, provided by Guzmán in [8]. Construction of completely separable mad families in $\omega]^{\omega}$ has been extensively studied in the past two decades. Furthermore, complete separable mad families in $\omega]^{\omega}$ have proved to be a powerful tool for constructing other set-theoretic objects. These include objects of interest in functional analysis, such as a masa in the Calkin algebra which does not lift to a masa of $B(H)$ in [21]. Shelah attained a breakthrough in [22], where he showed that a completely separable mad family exists assuming either $\mathfrak{c}<\aleph_{\omega}$, or there is no inner model with a measurable cardinal. However, the existence of a completely separable mad family in $\mathsf{ZFC}$ alone remains open. We shall show that the natural generalisation of a completely separable mad family to vector spaces (see Definition 5.5) exists in $\mathsf{ZFC}$.
@@
Our constructions for Theorem 1.5 and 1.6 utilise various results from the local Ramsey theory of block subspaces developed by Smythe in [23] and [24], along with a few additional results that we shall prove. The non-requirement of additional set-theoretic hypothesis is a result of the $\mathsf{ZFC}$ inequality $\operatorname{non}(\mathcal{M})\leq\mathfrak{a}_{\mathrm{vec},\mathbb{F}}$ by Smythe in [24]. Since $\mathfrak{s}\leq\operatorname{non}(\mathcal{M})$ holds in $\mathsf{ZFC}$, the vector space variant of the inequality $\mathfrak{s}\leq\mathfrak{a}$, i.e. $\mathfrak{s}\leq\mathfrak{a}_{\mathrm{vec},\mathbb{F}}$, is also a $\mathsf{ZFC}$ inequality. Readers may also refer to [16] and [17] for more constructions of mad families in $\omega]^{\omega}$ involving the cardinal invariant $\operatorname{non}(\mathcal{M})$.
The second part in this paper addresses the existence of full mad families of block subspaces over $\mathbb{F}_{2}$ in Solovay’s model. We utilise the abstract Mathias forcing developed by Di Prisco–Mijares–Nieto in [6] — the abstract Mathias forcing was developed to expand the local Ramsey theory of $\omega]^{\omega}$ to topological Ramsey spaces (i.e. closed triples $(\mathcal{R},\leq,r)$ satisfying the axioms A1-A4 introduced by Todorčević in [27]).
If $E$ is a countable vector space over $\mathbb{F}_{2}$, then the pigeonhole principle holds, so $(E^{[\infty]},\leq,r)$ is a topological Ramsey space. By identifying vectors in $E$ with their supports, this topological Ramsey space coincides with Hindman’s space $(\mathbf{FIN}^{[\infty]},\leq,r)$ studied in [3]. Filters in $\mathbf{FIN}^{[\infty]}$ have been studied by Blass in [3], and are known as ordered-union ultrafilters. Additional overlaps between Blass’ theory of ordered-union ultrafilters and Smythe’s theory of filters in [23] and [25] include the notion of stable ordered-union ultrafilters, which in Smythe’s terminology is known as a full (p)-filter. This allows us to study the local Ramsey theory of infinite block sequences in $E$ from the following perspectives:
  1. (1)
    Abstract local Ramsey theory by Di Prisco–Mijares–Nieto in [6].
  2. (2)
    Local Ramsey theory of block sequences by Smythe in [23] and [25].
  3. (3)
    The theory of (stable) ordered-union ultrafilters by Blass in [3].
In particular, the first two theories proposed versions of (semi)coideals and (ultra)filters in $E^{[\infty]}$, along with definitions of various selectivity properties. In §6 and §7, we shall show that these three theories coincide, so definitions may be used interchangeably and results proven in the individual theories may be freely applied to (semi)coideals and (ultra)filters in $E^{[\infty]}$.
Let $\operatorname{Coll}(\omega,<\!\kappa)$ be the Lévy collapse of uncountable cardinal $\kappa$ to $\omega_{1}$. Using the abstract Mathias forcing developed, we may apply the abstract Mathias forcing to obtain the following variant of a celebrated theorem of Mathias in [13]:
@@
A rough sketch of the proof of Theorem 1.7 (in the context of $\mathbf{FIN}$) was stated by Blass in [3], and he stated that it “will be shown, in a more general context” in [4], but the published version of [4] does not contain such a proof.
@@
Let $\kappa$ be a Mahlo cardinal, and let $G$ be a $\operatorname{Coll}(\omega,<\!\kappa)$-generic filter. Then there are no full mad families of subspaces over $\mathbb{F}_{2}$ in $\mathbf{L}(\mathbb{R})^{\mathbf{V}[G]}$.
@@
    We provide a concise introduction for various definitions and notations to block subspaces of $E$. We prove that finite intersections of block subspaces are also block subspaces (Lemma 2.3).
@@
    We introduce some definitions and basic results about almost disjoint families of subspaces. We show that $\mathfrak{a}_{\mathrm{vec},\mathbb{F}}^{*}$ is uncountable, addressing a remark by Smythe in [24]. We also show that the semicoideal $\mathcal{H}^{-}(\mathcal{A})$ is selective for any almost disjoint family $\mathcal{A}$ of subspaces — the non-requirement for $\mathcal{A}$ to be maximal is crucial in the proofs of Theorems 1.5 and 1.6.
@@
    We apply various lemmas and results proven in §3 and §4 to provide detailed proofs of Theorems 1.5 and 1.6.
@@
    We provide an overview of the axioms of topological Ramsey spaces A1-A4 introduced by Todorčević, and the abstract local Ramsey theory developed in [6]. We show that given a closed triple $(\mathcal{R},\leq,r)$ satisfying A1-A4, if $\mathcal{H}\subseteq\mathcal{R}$ is semiselective then $\mathbb{M}_{\mathcal{H}}$ satisfies the Mathias property and the Prikry property. This was proven in [6], but it was assumed that the almost reduction relation $\leq^{*}$ is transitive (which is not generally true, such as for the topological Ramsey space of strong subtrees). We provide a proof of this result without this assumption.
@@ 2. Vector spaces and block subspaces
We introduce some basic definitions and properties of vector spaces and block subspaces. Throughout this article, only non-zero vectors shall be considered (thus, for instance, for two subspaces $V,W$, we write $V\cap W=\emptyset$ if their intersection is the trivial subspace).
@@ 2.1. Setup and notations
Let $E$ be a vector space over a countable field $\mathbb{F}$, with a countable Hamel basis $(e_{n})_{n<\omega}$. For each $x\in E$, we may write $x:=\sum_{n<\omega}\lambda_{n}(x)e_{n}$, where $\lambda_{n}(x)\in\mathbb{F}$. This allows us to define the support of a vector $x\in E$ by:
@@
Note that $\operatorname{supp}(x)$ is finite for all $x\in E$, as $(e_{n})_{n<\omega}$ is a Hamel basis. This gives us a partial order << on $E$, where for any two vectors $x,y\in E$:
@@
Since such a block basis is unique up to scaling, we may conflate block subspaces of $E$ with infinite block sequences. This allows us to equip $E^{[\infty]}$ with a Polish topology by considering the first difference metric on infinite block sequences — that is, if $(x_{n})_{n<\omega}$ and $(y_{n})_{n<\omega}$ are infinite block sequences, then:
$d((x_{n})_{n<\omega},(y_{n})_{n<\omega})=2^{-\min\{n:x_{n}\neq y_{n}\}}.$
@@
  3. (3)
    If $A=(x_{n})_{n<\omega}\in E^{[\infty]}$ and $N<\omega$, we write $A/N:=(x_{n})_{n\geq N}$ [^1] This notation differs from that introduced in [25], where it was instead defined to be the tail sequence of $A$ consisting of vectors supported above $N$..
  4. (4)
@@
  5. (5)
    If $a=(x_{n})_{n<N},b=(y_{m})_{m<M}\in E^{[<\infty]}$, we write $a<b$ if $x_{N-1}<y_{0}$.
  6. (6)
    If $A\in E^{[\infty]}$ and $a\in E^{[<\infty]}$, we write $a<A$ if $a<r_{1}(A)$.
  7. (7)
    If $A\in E^{[\infty]}$ and $a\in E^{[<\infty]}$ we write $d_{A}(a):=\min\{N<\omega:a<A/N\}$.
  8. (8)
@@
  9. (9)
    If $Y\subseteq E$, we write:
    $\langle Y\rangle:=\operatorname{span}(Y).$
  10. (10)
@@
    $\langle(x_{n})_{n<\alpha}\rangle:=\langle x_{n}:n<\alpha\rangle.$
  11. (11)
@@
  13. (13)
    If $A\in E^{[\infty]}$, we write $E^{[\infty]}\mathord{\upharpoonright}A:=\{B\in E^{[\infty]}:B\leq A\}$.
  14. (14)
    If $A\in E^{[\infty]}$, we write $E^{[<\infty]}\mathord{\upharpoonright}A:=\{a\in E^{[<\infty]}:\langle a\rangle\subseteq\langle A\rangle\}$.
@@ 2.2. Intersection of block subspaces
Given $A,B\in E^{[\infty]}$, if $\langle A\rangle\cap\langle B\rangle$ is infinite-dimensional, then by the remark above it contains a block subspace. The following lemma shows that $\langle A\rangle\cap\langle B\rangle$ is itself a block subspace [^2] This lemma corrects footnote 8 of [24], which incorrectly stated that intersection of block subspaces need not be block.. This observation is central in ensuring that the proof of Theorem 1.5 and 1.6 do not require additional set-theoretic hypotheses.
@@
For an arbitrary $v\in E$ and $k\in\operatorname{supp}(v)$, recall that $\lambda_{k}(v)\in\mathbb{F}$ denotes the unique coefficient such that $v=u+\lambda_{k}(v)e_{k}+w$ for some vectors $u<e_{k}<w$ (or $u=0$ or $w=0$).
@@
$v_{0}$ $=\sum_{i\in\operatorname{supp}_{A}(v_{0})}\xi_{i}^{0}x_{i}=\sum_{j\in\operatorname{supp}_{B}(v_{0})}\mu_{j}^{0}y_{j},$ (2.1)
$v_{1}$ $=\sum_{i\in\operatorname{supp}_{A}(v_{1})}\xi_{i}^{1}x_{i}=\sum_{j\in\operatorname{supp}_{B}(v_{1})}\mu_{j}^{1}y_{j}.$ (2.2)
where $\xi_{i}^{0},\xi_{i}^{1},\xi_{j}^{0},\xi_{j}^{1}$ are all non-zero. We let:
$w:=\sum_{i\in\operatorname{supp}_{A}(v_{0})\cap\operatorname{supp}_{A}(v_{1})}\xi_{i}^{0}x_{i},$
@@
By the symmetry of the argument, it suffices to show that $\operatorname{supp}(w)\subseteq\operatorname{supp}(w^{\prime})$, and $\lambda_{k}(w)=\lambda_{k}(w^{\prime})$ for all $k\in\operatorname{supp}(w)$. Let $k\in\operatorname{supp}(w)$, so $k\in\operatorname{supp}(x_{i})$ for some $i\in\operatorname{supp}_{A}(v_{0})\cap\operatorname{supp}_{A}(v_{1})$. By the second equalities in (2.1) and (2.2), we have that $k\in\operatorname{supp}(y_{j})$ for some $j\in\operatorname{supp}_{B}(v_{0})$, and $k\in\operatorname{supp}(y_{j^{\prime}})$ for some $j^{\prime}\in\operatorname{supp}_{B}(v_{1})$. Since $B$ is a block sequence, $k\in\operatorname{supp}(y_{j})\cap\operatorname{supp}(y_{j^{\prime}})$ implies that $j=j^{\prime}$. Therefore, $j\in\operatorname{supp}_{B}(v_{0})\cap\operatorname{supp}_{B}(v_{1})$, so $k\in\operatorname{supp}(w^{\prime})$. By (2.1), we can conclude that:
$\lambda_{k}(w)=\xi_{i}^{0}\lambda_{k}(x_{i})=\lambda_{k}(v_{0})=\mu_{j}^{0}\lambda_{k}(y_{j})=\lambda_{k}(w^{\prime}),$
@@
$z_{m}$ $=\sum_{i\in\operatorname{supp}_{A}(z_{m})}\xi_{i}x_{i}=\sum_{j\in\operatorname{supp}_{B}(z_{m})}\mu_{j}y_{j},$
where $\xi_{i},\mu_{j}$ are non-zero. Let $i_{0}:=\max\{i\in\operatorname{supp}_{A}(z_{m}):i<n\}$ and $i_{1}:=\min\{i\in\operatorname{supp}_{A}(z_{m}):i>n\}$, and let $k_{0}:=\max(\operatorname{supp}(x_{i_{0}}))$ and $k_{1}:=\min(\operatorname{supp}(x_{i_{1}}))$. We define:
$w:=\sum_{\begin{subarray}{c}i\in\operatorname{supp}_{A}(z_{m})\\ \max(\operatorname{supp}(x_{i}))\leq k_{0}\end{subarray}}\xi_{i}x_{i}.$
@@
Observe that for each $j\in\operatorname{supp}_{B}(z_{m})$, either $k_{0}<\min(\operatorname{supp}(y_{j}))$ or $\max(\operatorname{supp}(y_{j}))<k_{1}$ — otherwise, if $j\in\operatorname{supp}_{B}(z_{m})$ is such that $\min(\operatorname{supp}(y_{j}))\leq k_{0}<k_{1}\leq\max(\operatorname{supp}(y_{j}))$, then by the block sequence property of $B$, we have that for all $j^{\prime}\neq j$, either $\max(\operatorname{supp}(y_{j^{\prime}}))<k_{0}$ or $k_{1}<\min(\operatorname{supp}(y_{j^{\prime}}))$. Since $k_{0}<\min(\operatorname{supp}(x_{n}))\leq\max(\operatorname{supp}(x_{n}))<k_{1}$ and $j\notin\operatorname{supp}_{B}(z_{n})$, this implies that:
$\operatorname{supp}(x_{n})\cap\bigcup_{j^{\prime}\in\operatorname{supp}_{B}(z_{n})}\operatorname{supp}(y_{j^{\prime}})=\emptyset,$
@@
Note that for all $k\in\operatorname{supp}(z_{m})$, either $k\leq k_{0}$ or $k_{1}\leq k$. Thus, for all $j\in\operatorname{supp}_{B}(z_{m})$, since $\operatorname{supp}(y_{j})\subseteq\operatorname{supp}(z_{m})$, if $\max(\operatorname{supp}(y_{j}))<k_{1}$ then $\max(\operatorname{supp}(y_{j}))\leq k_{0}$. Similarly, if $k_{0}<\min(\operatorname{supp}(y_{j}))$ then $k_{1}\leq\min(\operatorname{supp}(y_{j}))$. Therefore, we have that:
@@
then we have that $\operatorname{supp}(w)=\operatorname{supp}(w^{\prime})$. Furthermore, for any $k\in\operatorname{supp}(w)$, $\lambda_{k}(w)=\lambda_{k}(z_{m})=\lambda_{k}(w^{\prime})$, so $w=w^{\prime}$. It’s clear from the definition that $w\in\langle A\rangle$ and $w^{\prime}\in\langle B\rangle$, so we have that $w\in\langle A\rangle\cap\langle B\rangle$, as desired. ∎